OpenStudy (anonymous):

Evaluate (y – 3) ÷ x when y = 15 and x = 3.

OpenStudy (anonymous):

plug in 15 for y and 3 for x.

OpenStudy (anonymous):

(15-3) / 3 = 12 / 3 = you got this

OpenStudy (anonymous):

12/3 = 12/3 4=4